OpenStudy (anonymous):

I have to add the following z/z^2-49 + 2/z+7

OpenStudy (anonymous):

\[\frac{z}{z^2}-49+\frac{2}{z}+7=\frac{1}{z}-49+\frac{2}{z}+7=\frac{3}{z}-42\]

OpenStudy (anonymous):

actually I should of wrote it out as z/(z^2-49) + 2/(z+7), does that make the answer different??

OpenStudy (anonymous):

yes, it does

OpenStudy (anonymous):

start w a common denom of (z+7) (z-7)

OpenStudy (anonymous):

thats about as far as I got!!

OpenStudy (anonymous):

z+2(z-7) all over (z+7)(z-7)

OpenStudy (anonymous):

then the z-7 cancels out, right??

OpenStudy (anonymous):

z+2z-7 -------- z^2-49

OpenStudy (anonymous):

So is that the answer??